12 Mar 2008: There was colliding Au beam at 9 GeV/n in both PHENIX and STAR last night. STAR received their goal of 5000 good events. PHENIX was also satisfied with the beam delivered to them. By 10:30PM Todd was beginning to tune up 5 GeV/n Au beam. Au beam was brought to AGS Extraction. At 02:30 there was a Quench Link Interlock in b4-dh0 which resisted attempts by CAS to fix it. Finally a card was replaced to allow Blue to ramp without b4-dh0 (which unfortunately shares a power supply with the DX magnet required for Yellow Beam) around 06:00AM. By 10:00AM Au beam at 5 GeV/n had been sent half way around the Yellow ring. There was no morning meeting this morning to allow Todd to continue Low Energy development.
11 Mar 2008: Good progress getting low
energy Au colliding. STAR saw real vertices for central and peripheral
collisions at 9GeV/n root(s). They even managed a store of 1/2 to 1 hour
length, although the collision rates are low at the end of such a store.
Difficulties reading the PHENIX scalers have made it impossible to verify
Au-Au collisions at PHENIX.
Todd determined that running with h=366 does not allow collisions at both
STAR and PHENIX simultaneously. He advocates moving to h=369 (and from E=9
to E=8 GeV/n) which will allow collision at both. This requires retuning
injection, taking 2-4 hours. An alternative would be to shift to 5 GeV/n.
STAR needs a vernier scan, as does PHENIX. Measured emittances out
of the AGS and in RHIC were 3pi in both.
There was a discussion of the best plan for the day. Current stores are
20-30 minutes in length. We will work on developing collisions in
PHENIX, Once there are verified collisions, they will turn on the PHENIX magnet
and take some luminosity. Then RHIC cogging will be adjusted to provide
collisions at
STAR. STAR will take 5000 events. STAR will perform a vernier scan.
During this time PHENIX will be studying their initial data. When STAR has
completed their 9 GeV dataset, PHENIX will get collisions until their dataset is
also complete. Then there will be a vernier scan at PHENIX. When
that is complete, at approximately 5PM the switch to 5GeV/n will begin.
Injection will have to be retuned. Then an attempt will be made to
circulate 5 GeV beam in RHIC. Once there is beam in RHIC, there will be a
study of Inter Beam Scattering effects by looking at the beam lifetime at a
range of intensities. This completes the RHIC program.
10 Mar 2008: There was work done to locate
the Blue and Yellow beams at the JET so that the polarization of both beams
could be measured at the same time. This was accomplished on Saturday,
with Blue at (0, -1)mm and Yellow at (0, -7)mm.
Chromaticity adjustment for the ramp in Yellow. This did not make the
polarization either better or worse.
Fill 10000 had intensities of 1.7 and 1.3e11 giving 9kHz at PHENIX ZDC.
Polarization was 53/51 (injection) dropping to 48/44 (Store)in Blue/Yellow.
The beam separation at the JET was decreased from 6 to 5.5mm. The end of
fill 10000 was devoted to a vernier scan for PHENIX.
Most of last night was spent trying to get injection working in Yellow.
Five or six stores were lost for a variety of reasons: the permit was pulled by
the NMC at PP2PP, the permit got pulled by the Snake, and the RHIC RF power
supply BS2 tripped off. In order to get a store in for PP2PP testing we
had to disable the NMC interlock, leave the RF Station BS2 off, and fill RHIC
with only 1e11 in both rings.
PHENIX has turned their magnets off. STAR will leave their magnets on.
PHENIX requires an access at the end of polarized proton running. STAR
also requires a 30 minute access. Paul Sampson has requested that the two
experimental accesses be scheduled sequentially to minimize the number of gate
watch personnel needed.
Todd will work to establish root(S)=9 GeV Au beam in both Blue and Yellow.
Set-up using beam energies of 4.592, 2.6 and 2.5 GeV per beam total energy.
This requires flipping to polarity of the sextupoles, expected to take ~3 hours
beginning at 8:00AM. Au beam has been delivered through the Booster
already. They are running in PPM so that NSRL can run simultaneously.
Anatoli has requested 24 hours of continued polarized proton development work in
the LINAC to help understand the source polarization. This impacts the
duty factor for BLIP, reducing the beam delivered to BLIP by ~15%.
NSRL is running Carbon at 209 and 1000 MeV/n today, in PPM with Au.
07 Mar 2008: Cryo test yesterday we lowered the
flow rate of liquid helium to observe the 10 Hz orbit oscillations.
Results are not conclusive. It appears that the oscillation frequency is
changed, perhaps to random, but the amplitude of the oscillation is about the
same. The cryo test was performed by reducing the flow rate to 100 g/s,
which gives a temperature increase between recoolers of 100mK. The flow
rate can be dropped to 60 g/s with a delta-T of 190mK. The 10Hz study
could be repeated at the 60 g/s flow rate to look for more definitive results.
There will be an attempt to use the JET to measure both Blue and Yellow
simultaneously.
Fill 9992 had polarization of 56/53 at injection, dropping to 56/45 at store.
At lower intensity (1.5/1.4e11) the Yellow ramp efficiency has recovered to 98%.
Carbon Polarimeter measurement seems to fail frequently. There needs to be
a new set of instructions, and improved reliability in the Controls Software.
Mei Bai is working with Todd to perform a polarization measurement on the Ramp.
Waldo will continue to do snake scans.
PHENIX is interested in two vernier scans; one long at the end of a store.
06 March 2008: APEX ran yesterday from 9:00 -
16:00. Measured spin tune at injection.
There was a measurement of Yellow polarization at injection and store (fill
9988) that appeared to show improved polarization transmission on the ramp.
For fill 9989, we filled 1.7/1.6e11 for 13kHz in PHENIX ZDC at store, with large
Yellow vertical emittance initially. Polarization went from 52/38 to 49/40
at store. Yellow ramp efficiency is only 89%.
PHENIX made a brief access to attempt to fix a Drift Chamber HV problem.
It was not successful.
This morning PHENIX made another access to address the DC HV problem again and
determined they need a longer ~1 hour access to replace a HV mainframe.
At the end of the first store today there will be a test of the 10Hz
oscillations by turning off the cryo flow. Todd will coordinate these
tests.
Due to problems with the AGS Corrector magnet, beam is down this morning, and
both STAR(PP2PP) and PHENIX will be taking accesses of ~1 hour to address
problems.
05 March 2008: Nonlinear chromaticity
corrections yesterday. Not much improvement.
Running at high per bunch intensities, 1.8e11 with polarizations of 50/45 at
store. 55/50 at injection. PHENIX ZDC rates were as high as 14.5kHz.
Also, in the middle of the night, the PHENIX ZDC luminosity increased by 15% for
no apparent reason. STAR ZDC may have seen a slight increase at the same
time, but not in the BBC.
There was a cryo lead flow fault that kept RHIC power supplies interlocked off
for several hours.
It was discovered that the D1 extraction bump used for NSRL disrupts the
polarized proton delivery to the AGS. It was turned off without any
apparent effect on NSRL.
Today there will be APEX studies all day, 9:00 - 16:00. The end of APEX
will be devoted to looking at the effect of cryo pumping on the 10Hz orbit
oscillations. The cryo will be turned off at 3:00PM, and after ~45 minutes
to let the system equilibrate, a measurement of the orbit will be taken with
BPMs to see the strength of the 10Hz oscillations. When the measurements
are completed, the cryo will be turned back on. It is hoped this
measurement will provide input to the new working point problem.
04 March 2008: We had several accesses during the afternoon yesterday. The first was for STAR to address their DAQ problem. Initial suspicion of high radiation levels turned out to be unfounded. DAQ instability was due to a misbehaving board. This took several attempts to correct. Another 30 minute access was made to unclog the JET Hydrogen nozzle. A short access was made to fix the Blue beam synch problem in the 7C alcove.
03 March 2008: NSRL started running experiments this
morning. No issues.
In RHIC, the Beam intensity is up. Polarization down.
Did a scan in the AGS of intensity vs polarization, and it was not possible to
get more than 50% at these intensities. 53/47 at injection, down to 51/38
at store. Yesterday's beams were 1.48/1.43e11 with 11.5kHz PHENIX ZDCs.
This morning we ran 1.61/1.56e11 at 12.5kHz, decaying within 15 minutes to
11kHz. Polarization is 54/53% at injection and 53/40% at store. The
last two stores appear to show the polarization in Yellow is no longer dropping
throughout the store. There is no known change in configuration.
STAR reports that the electronics on the front of the TPC is failing, due
probably to high radiation levels from these high beam intensities.
Today, after the morning fill we will swap the orbit of the Jet from Blue to
Yellow. Yun will work on the nonlinear chromaticity correction. This will
take the beam for ~2 hours. The Blue beam synch was lost in the morning
store, so there will be a short access to allow this to be reset.
JET requires a half hour access for service. Before the access, the JET requires
a 4 hour warm up, then access, and then a 4 hour cool down. Currently the
JET is taking off-axis data, and they do not want to do the servicing until
after this fill, and probably not before tomorrow.
BLIP is going to 24/7 operation. Peter Ingrassia will ensure there is
operator/CAS coverage..
29 Feb 2008: Returned to low beta* set-up
last night, after a one hour access for PHENIX to "flip a switch" and STAR to
search for and fix instabilities in their data acquisition. With
56x56, the ZDC rates at PHENIX were 8kHz, very good. But lifetime was very
poor. An attempt was made to fill with 109 bunches at 1e11, but the BLAM
shut the work down for an hour, and the decision was made to return to pp83
(rather than pp83lowbeta) ramp file for the night's physics store at 1.3e11 with
105x98 bunches giving 9kHz for PHENIX ZDC. Polarization was 48/40 at
store, with 54% from AGS, running with Run08 set-up. Haixin is working on
transitioning to Run06 settings. After spending the night working on it,
we are probably ready to inject with emittances and polarization at the level of
Run06.
After the morning store is dumped, we will swap the JET from Yellow to Blue.
There will also be work on tune adjustment to get the tunes on opposite sides of
the diagonal.
28 Feb 2008: Maintenance Day yesterday went
well. PHENIX got their MuTR trigger work done. Most of the PP2PP
development work got done. Booster and AGS work got done, including the
target replacement for the AGS Polarimeter and the polarized proton source
maintenance. Recovery from Maintenance also went well, with a 12 bunch
store and a physics store by 8:00PM.
The APEX work on beta*=60cm development was stymied by loading two incorrect
ramp files. pp83 instead of pp83lowbeta. They need a couple of good
ramps with the correct ramp file in order for collimation work and tuning to be
completed.
The STAR ZDC rates are higher than the PHENIX ZDC rates now. This is a
change for this run, and it is not understood . Wolfram is concerned that
we are using the wrong cross sections for calculating emittances. Previous
running used 41 and 50 nb for PHENIX and STAR, whereas this year it appears to
be 31/29 nb. We need to use all our tools to study the emittance and come
to a consensus.
STAR requests an half hour access to work on their data acquisition system to
improve stability. This will occur immediately after the beam is dumped at
13:00. Prior to dumping the beam, there will be a polarization measurement
and a 10 minute study for the stochastic cooling.
26 Feb 2008: Attempts to reproduce the problem
with the Blue abort kicker did not succeed. There is currently no evidence
for mistiming of the Blue abort kicker. At this point, there is no good
evidence that there was ever a problem with the abort kicker timing.
There was a short access yesterday to unblock a cooling water pump for the
holding magnet in the JET. PHENIX took the opportunity to swap out a LV
Power Supply. STAR/PP2PP also made an access to exchange a crate.
Polarization has been poor most of the night; <40%. Some of this is due to
reduced polarization out of the AGS which was down to 45% for at least one fill.
Today, immediately after the 9:00AM meeting we will swap the orbits for the JET.
We will attempt to revert to the Run06 settings in AGS. This may improve
emittance blow-up, and improve polarization. There will be work on
polarization today.
APEX will begin tonight at 8:00PM til 8:00AM Wednesday morning when the
Maintenance Day will begin. The primary task of APEX will be to revisit
the beta* squeeze.
Following the Maintenance Day on Thursday we will try to complete the nonlinear
chromaticity correction.
25 Feb 2008: Power dip on Sunday brought down
the fill at 10:36AM, taking most of the day to recover. Beam was not
brought back in RHIC until after midnight this morning. Blue abort kicker
timing was wrong, late by 200ns, which made for dirty dumps. To compensate
for this, the first 4 bunches are filled only to low intensity.
On Saturday night we went back to measuring the JET polarization in the Blue
ring at about 9:00PM. On Sunday the JET holding field magnet tripped off
(again). No JET measurements can be made until they get a short access to
reset the cooling loop fault.
Polarization is low in both rings, <40% at store. Blue Vertical Target #4 broke,
so we had to switch to the thick Blue Vertical target #5 for polarization
measurements. On Friday we lost Yellow Vertical Target #1. We are
losing one target a week this year. Last year we were losing one target
every two weeks. Statistics are poor, but things do not look good for the
remaining two weeks of the run.
Today there will be work on fixing the timing of the Blue abort kicker.
And continuing attempts to improve the polarization and intensity.
21 Feb 2008: APEX running went well yesterday.
Work on the 60cm b* was the focus. There was
additional work to study the Polarimeter. measuring the polarization profile..
Recovery from the APEX time went quickly. There was a 6 bunch store
brought up first to check the ramp, and this store was used to switch the JET
from the Blue ring to the Yellow. The machine went back to 100cm
b* for production running through the night during
which we put up two good stores.
At the end of the morning store we will take half an hour to study polarization
profile and emittance growth in both Blue and Yellow.
Today we will try to finish the 60cm
b* development. We will also spend some time
offline developing the ramp polarization measurement.
JET requires that the Blue beam be moved to a new location where backgrounds
should be better. The measurement takes about 2 hours. This will be
done with the next good store.
20 Feb 2008: Nine hour store yesterday with
polarization 55%/48% in B/Y. Vernier scan completed, manual cogging tested
and it was found that Blue bucket 1 (bunch 1) was colliding with Yellow bucket 4
(bunch 2) at PHENIX during fill 9798. Two ramps failed due to no
separation at JET. The second pulled the permit. Reverted to an old
fill. 1.3x1.3e11 with 109x109. PHENIX ZDC was ~10kHz at start of
fill. PASS dropped the next store after 4 hours. PASS tripped
another gate, like last week.
APEX is running all day. Preliminary results are that the 60cm
b*-squeeze is working well, so previous plans to
take 4 hours tomorrow to try new ramps at intermediate b*.
will not happen. We will go straight to 60cm. This will be attempted
tomorrow during the day. At the end of APEX today we will switch the JET
from Blue to Yellow.
19 Feb 2008: Over the long weekend there were
problems with the microswitches on the door at 4 o'clock. The PASS system brought
down RHIC on Sunday at 3:11AM, and initial work on the PLC was not successful.
Replacing the switch did not help the first time, since the replacement switch
was bad. Replacing the switch again solved the problem, but the PASS
system brought RHIC off for over 12 hours. Very frustrating.
The timing of the injection kickers was found to be wrong. Fixing this
reduced the emittances in 109 bunch running.
Development of the 80cm b* is expected to
continue today.
The JET has accumulated ~1 million events in Blue, and expect to move to
Yellow. This should give a polarization measurement at the 8% statistical
level.
STAR says Blue and Yellow polarizations are about the same. The STAR
local Polarimeters indicate polarization that is consistent with ~50%.
Source polarization is ~80%. Polarization out of the AGS at Flattop is
~60%. We have the tools to measure the polarization profile in both
horizontal and vertical in the AGS. We should do this now.
15 Feb 2008: APEX running from 5:00AM -
5:00PM yesterday. First test ramp after APEX failed due to leftover
chromaticity settings. Dejan remembers this occurring "many times" in the
past when recovering from APEX. This is one of many instances of loss of
configuration control.
Finally there was an overnight store with 0.8e11 in 109 bunches giving 2.7kHz at
PHENIX.
The tape did not remove the separation bumps. Activating the ramp in the
RampEditor did.
Blue polarization is stable near 50% throughout the store. Yellow
polarization is lower, and drops throughout the store. There is no
understanding of why the Yellow polarization drops.
Wolfram emphasized the importance of performing an emittance measurement with
the Polarimeter at the beginning of the store.
Blue beam decay is high (110% per hour) at beginning of the store. The
Blue emittance at beginning of store is also large (44pi), explaining the high
losses and short lifetime. Lowering the Blue horizontal tune by 0.001
seems to help somewhat. Blue bunch length shrinks initially, then starts
to grow after one hour.
Tunes are on opposite sides of the diagonal with Blue below and Yellow above.
Blue should be "happier" than Yellow, but it isn't.
Today we will change the ramp tunes to separate the two beams and prevent the
emittance blow-up.
Run 109 bunches at 0.9e11 evening and owl shift.
Perform Polarimeter Target Scan at beginning and end of each store, and measure
polarization every two hours. Call experiments before scans.
Measurements of Blue beam emittance at the end of the store today show values of
~60pi with Target #1 (thin) but only 21pi with Target #5 (thick). This
suggests that the large emittance observed in Blue is a reflection of target
peculiarities. Meanwhile, ZDC-based emittance points at higher values,
30-40pi. In the process of exploring emittance measurements, it was
discovered that all four of the Yellow Horizontal Targets has been lost.
We have no method of measuring the Yellow Vertical Polarization Profile any
longer.
The JET lost an RF Cooling coil today and need a 20 minute access today before
physics stores begin. PP2PP also needs a 10 minute access that can fit
behind the JET work. Both of these tasks can probably fit behind a rotator
ramp (~30 minutes).
14 Feb 2008:
13 Feb 2008:
12 Feb 2008: AGS J-10 Dump Bump Power Supply had over
temperature alarm, requiring flushing of the cooling lines.
Bad settings from the collimators were discovered (leftover from d-Au
running) due to high background rates at both PHENIX and STAR. A new set
of collimator settings was loaded giving good background rates in both Blue and
Yellow for both PHENIX and STAR.
Much of yesterday was spent working on the rotator ramp. In addition
the number of bunches was increased from 56 to 109. On the first attempt
at 109x109 with 1.1e11 per bunch, snake losses early in the ramp pulled the
permit. A second store with 0.65e11 was lost at store. So a decision
was made to run 56x56 at 0.6e11 overnight. Polarimeters measured more
68/58% polarization in the Blue/Yellow. Many problems having to do with
hanging the Polarimeter FEC, and operator inexperience made for rough going.
PHENIX wants to take a store to measure the polarization state now that we
have gone to radial polarization at 8 oclock. During this store, PHENIX
will commission their asymmetry detectors. Comparing to CNI detectors
requires a polarization measurement every hour.
11 Feb 2008: BBQ blew up emittances. 15
minutes with BBQ on caused emittance to blow up from 13pi to 28pi, due to
increased voltages for BTF.
Most of the day yesterday was spent working on rotator ramp for pprot83.
Power dip brought down the AGS which gave trouble most of the night.
Provided collisions overnight without rotators, although intensities were low
0.7-0.8e11 in both.
Two Quench Link Interlocks overnight. One was due to a fibre
optic interface card failure, and the other was due to a Yellow Main PS.
JET is still down due to a failure of the Breit-Rabi polarimeter gauge.
Need 4 hour access.
So today we need to continue rotator ramp development. Nonlinear
chromaticity correction.
10 Feb 2008: JET had an access to fix a water
circulator issue. Operational by 4PM. During the JET access we found
Yellow CNI bias voltage at 15 volts instead of 90. Raising the voltage to
90 increased the measured polarization from 30 to 50%. Today we need to
understand what is needed to fix the Breit-Rabi Polarimeter for the JET.
The JET also requires a pump replacement and an ion gauge replacement.
These items will require an access of ~8 hours. During the first good
store after the access yesterday, we took a minimum LOTO access to repair a
Booster RF cavity.
We ran a 1.1E11 56x56 bunch store last night, with vertical polarization for
PHENIX. They have now finished calibrating there polarization detectors,
and are ready to go to Radial Polarization,.
Both experiments report good background rates at these operating conditions.
Calculated emittances from ZDC rates are >30pi. IPMs show only 20pi.
Today we need to understand the emittances, perform rotator ramp development and
collision tuning.
09 Feb 2008: Access for STAR and PHENIX went
well, although the PHENIX Iris scanner failed again. The bo6-qd3-ps power
supply was swapped during the access.
Yellow feedback ramp worked with efficiency of 99.5%. Lowered cavity
voltage from 100 to 75kV and Landau phase was adjusted.
LINAC module 6 was arcing last night and had to be fixed. There was work
on collimation during the LINAC repair.
The Yellow polarization was zero percent after feedback ramp. Work on
Yellow polarization adjusting tunes and chromaticities. Measured 30% at
store.. Filled with 56x56 store with 1e11 showing poor emittance and
lifetime. Both improve after tune change. There was a Blue quench at
4:30AM caused by running the Blue Polarimeter target frame into the beam.
This was accomplished by people using the pet pages to move the targets instead
of the Polarimeter GUI. The users should NEVER use the pet pages to move
the targets. Previously when the beam hit the target frame it destroyed
all the targets on that frame. .
The main activity for today is to improve the Yellow polarization. Te
question is how can the Blue polarization be 45% in Blue and only 30% in Yellow
at injection? Polarization extracted from AGS is identical by definition.
All the transfer elements in the AtR that effect the polarization are common to
Blue and Yellow. So how can the polarization at injection be so different?
Could it be a problem with the calibration of the Polarimeters? If
Polarimeter is ready, measure polarization on the ramp.
Check Yellow snakes if Yellow and Blue polarization are different.
08 Feb 2008: LINAC Failure lost 2 hours.
Blue stochastic cooling test, moved the storage and Landau RF phases at store
for Blue.
Yellow RF had wrong gamma_t value in DSP code.
Necessitated half bucket timing shift at store (17ns) and wrong
chromaticity. DX BPMs at store still
need to be adjusted.
Non-linear IR correction needs work.
Overnight collisions with 36x36 1e11.
PHENIX ZDC of 2kHz, 50kHz in BBC.
Today, continue Nonlinear IR correction in Blue (Fulvia), Nonlinear chromaticity
work (Yun), Feedback ramps and decoupling (Yun), Ramp tuning for polarization,
keep tune swing. Overnight
collisions at 1.1e11 56x56. With pp83 ramp, not pp83ro.
I.e. continue with vertical polarization until PHENIX gets a good run at
40% polarization.
CNI measurements every hour.
Get CNI into CDEV.
Continue Machine Development over the weekend.
STAR needs a half hour access to fix a trigger.
PHENIX will take advantage of the access to replace a LV PS.
PP2PP will also get work done inside.
This access will take place as soon as the morning meeting is over.
STAR reports a twenty second ripple in the background from Yellow.
07 Feb 2008: Spin pattern checked at Fill and it is OK
Measured large emittances (50pi) due to quad polarity reversal in AtR.
Fixed.
Feedback ramps got to 99% efficiency.
Nonlinear chromaticity correction
Attempted feedback ramps using store tunes.
Failed twice and abandoned.
Back to tune swing ramps.
Measured polarization at store. 36%
in Blue,10% in Yellow.
Rotator ramp set up during evening shift.
Second rotator ramp gave 97% blue and 89% yellow.
Blue horizontal store emittance at 20pi.
Wrong Blue spin pattern.
Provided collisions over night with 37x37.
7.5kHz at PHENIX, 120kHz BBC at both experiments.
JET set-up during store.
Tested 10Hz IR feedbacks. BPM
signals noisy; setpoints saturate.
Yellow horizontal store emittance 32pi, depolarized.
Next store gave 18pi with 15% polarization.
Blue decay still shows one minute periodicity.
Continue ramp tuning including polarization..
Do feedback ramps, nonlinear chromaticity correction.
IR nonlinear correction.
Check 10Hz feedback system..
Increase intensity and number of bunches.
There is an impending LINAC failure that needs a 1-2 hour access behind a store.
STAR requests an access for ~30 minutes to fix a stuck trigger bit.
Standing request for access by PP2PP.
PHENIX has requested a one hour access some time this week in order to
fix a TDC.
PHENIX needs to make measurements with vertical polarization to calibrate their
detectors. Then another ramp with
radial polarization for the remainder of the run.
Rob Michnoff reports that he observed the same kind of timing shifts between
d-Au and pp (ie 1ns at PHENIX and 17ns at STAR) and he suspects that there was a
phase shift in the V124 module providing the timing pulse.
06 Feb
2008:
After abandoning the new working point, reloaded the old magnet settings from
Run06. Circulating beam in Blue and Yellow by 5:30, with a first ramp by
7:10. Store dispersion looked good by 4:30 AM. A 12 bunch ramp was
performed with cogging in an attempt to see collisions. There were no good
collision signals yet. We will try for collisions again tonight.
An AGS Cooling failure took the beam from 6:30 - 9:00.
Polarimeter measures emittances of 50p in both
rings.
05 Feb 2008:
Experiments got collisions last night.
STAR saw a global timing shift of 17 ns with lots of background. PHENIX
saw a 1ns shift, and 15kHz of Beam-Beam counters coincidence counts. STAR
saw 37kHz with spikes 3 times per minutes where the instantaneous rate from the
Blue beam jump up by a factor of 50. This correlates with a Beam Loss of
200% per hour. Yellow beam looks great. Blue beam is very touchy,
shows a 1 minute oscillation in Wall Current Monitor readings. Haixin made
a measurement of the Blue beam emittance and saw 27pi mm-mr. This is quite
large, but agrees with emittance values derived from the STAR event rate.
Lowering the tunes by 0.02 improves the Blue orbit a lot (Todd) but we don't
know what effect this might have on polarization. We need to put the Blue
Vertical Damper back in.
Last night, just as RHIC was ready to give beam to the experiments, the PHENIX
Pad Chamber HV power supplies tripped off. They will need an access to fix
the problem.
There was a quench link interlock at about 9:00AM this morning. This
appears to be due to a power supply failure. As experts go in to diagnose,
PHENIX will take the opportunity to take an access to work on their Pad Chamber
HV power supplies.
Today we will explore tunes around 0.91.
Repair Yellow BBQ. Do feedback ramp.
Ramp tuning for transmission, store lifetime, polarization.
Polarization measurements.
Rotator ramp development (Waldo) unless we change the Blue ramp in which case
the rotator ramp would be postponed until after the Blue ramp is completed and
retuned.
04 Feb 2008: Beams are now in the center of the beam pipe,
after RF adjustments of ~40Hz corrected for radii that were off by ~1mm.
Number of bunches increased to 37 with Ramp efficiency up to 90%. There
were some first collisions overnight. Blue tunes are at 0.93/0.94 (instead
of 0.94/0.93) Blue beam loss shoot up every minute or so.
There was a snake quench this morning. It occurred when the loss monitors
showed a hit at the snake, even through the AtR switching magnet was off.
There was a discussion of how this could happen, or even whether it could
happen.
STAR does not have its ZDCs turned on, so we can't see collisions at STAR.
For today, machine development is to swap the Blue tunes. Todd should look
into the Yellow optics. Vadim will continue working on orbits at the IPs, and
separation bumps.
This evening, we should measure Blue profile with Polarimeter; verify injection
matching improvement. Verify the fill pattern at store. Measure
polarization at store. Haixin reports that currently the polarization at
extraction from the AGS is 60% at 1E11, dropping to 50% at 1.5E11.
03 Feb 2008: AGS problems kept beam off until 1:00PM
yesterday. PHENIX made an access (after troubles with Iris scanner)
Much work on improving orbits and tunes. Transmission efficiency up to 81%
in Blue and 54% in Yellow.
Continuing ramp development. Implement store tunes (Blue 0.96/0.95 Yellow
0.695/0.685). Measure Blue profile with Polarimeter and verify injection
matching improvement, and fill pattern.
02 Feb 2008:
Call in number for the bridge line is 631-344-8261, for both Saturday and
Sunday.
Blue beam made it to store (with feedback off). Ramp efficiency was 72%
with ARTUS on. Blue store orbit was 2mm RMS after first ramp. Micado
corrections brought it down to 0.4mm. Blue store tunes were 0.92/0.93 with
excellent lifetimes. Unfortunately the AGS RF Station B is down, and will
take a few hours to fix.
Yellow is still troublesome. We have to get Yellow tune measurements to
work, with BBQ and ARTUS. Then continue ramp developemtn on both rings.
Once we are happy with the ramp, we will have to measure the Blue profile
with polarimeter. Verify injection matching improvement. And verify the
spin pattern in "Fill" mode.
Tomorrow Waldo is on shift during the day, and he can work on rotator ramp.
01 Feb 2008: Ring-to-Ring synchro work took most of the day yesterday. So Ramp development was delayed until today. It was discovered that the Chromaticity in Yellow was off by one unit.
31 Jan 2008: Slow Factor 1 set up for ramp up, but
strill at Slow Factor 3 for ramp down.
Ring-to-Ring synchro Set Up
Polarimeter Set-Up and polarization measured at injection. Yellow is at
30%, Blue at 20%. Spin pattern in toggle mode is OK. Need to check
Fill Mode.
Ramp development using tune feedback today.
STAR requested an access to work on their trigger today.
30 Jan 2008: It took much longer to set up the
Snakes yesterday. RF Capture was established in both rings, but without
ring-to-ring synchro yet.
BPMs were timed in with fixed timing, ARTUS was set up, but we still need
time for BBQ and Polarimeter Set-Up.
RHIC Chromaticity failed in Yellow. Tunes set to nominal ramp values.
Tested store tunes in Blue. Both rings manually decoupled.
Both injection efficiency and lifetimes are excellent.
Blue vertical injection damper has no effect. Blue orbit corrections
need to be scaled down by a factor 2. Dipole trims have opposite effect
than in run 6. Blue tq polarities are OK. Blue ARTUS kicks cause
visible beam loss. May need to reduce strength..
Ramp work on pp80 for slow factors 1 (up) and 2 (down) did not get started
yesterday, so it is scheduled for immediately after the morning meeting.
PP2PP will work behind this store.
When beam comes back, we will finish setting up ring-to-ring synchro, and finish
setting up the instrumentation BBQ and Polarimeter.
There will be work today with Rob on re-checking dampers, Yun will work on
decoupling, Todd will change BPM to Gated Timing. Late today we will begin
ramp development using tune feedback.
29 Jan 2008: Maintenance/Set-Up for pp started at 8:00AM
yesterday. Power supply work went well, completing it in one "day".
Beam was brought to RHIC after midnight. There was circulating beam in
Blue and Yellow with good lifetime.
Today there will be work on lifetime, RF capture, and Instrumentation.
Beam out of the AGS has been 1.7E11, with a polarization at 1.5E11 of 40%.
STAR (pp2pp) has requested a 2 hour access during the Power Supply ramp
development work for Don some time today. Don's work will take ~4 hours.
We need to bring on the Snakes in RHIC before RF capture. Snake work
will be scheduled first, from 9:00-12:00. RF Capture will follow from
12:00-16:00. Power Supply ramp development will follow 16:00-20:00.
28 Jan 2008: Deuteron-Gold running ended this morning
at 8:00AM after 12 hours of APEX. Switch-over to polarized proton
running is underway. Beam is expected in RHIC tonight. If it is not
possible to complete the power supply work today, there will be limited work in
the AGS to improve polarization through the night, and finish the RHIC PS work
on Tuesday. The main tasks include switching the polarity of Q7 and the sextupoles..
PHENIX had problems over the weekend with a failure of a BBC temperature
interlock that disabled all of their triggers. A first attempt to reset
the crate mid-day Saturday did not cure the problem. A subsequent access
during which time the BBC temperature module was replaced finally brought PHENIX
back on-line.
Preparation for pp running includes bringing protons through the AGS at full
energy. Polarization of >50% has been measured in the AGS. But no
beam has been brought out of the AGS.
25 Jan 2008: Many difficulties made for inconsistent running This included a PLC failure for the 10 o'clock cryo. A Yellow loss monitor that was probably real, and a Blue loss monitor that dropped 4 stores before the troublesome loss monitor was masked out of the interlock. At the same time, the Blue intensity was reduced to allow for a store to be completed. This finally got the beam back, without a fundamental understanding of the source of the problem. Blue intensity was brought back up in the next stores, so it was most likely a loss monitor problem that caused the dumps.
24 Jan 2008: At the end of APEX, STAR requested an
access to allow them to cycle the power on the Forward Meson Spectrometer (FMS)
trigger system and Muon Telescope. A Blue Quad tripped on the first ramp
after the access, but recovery from APEX was relatively quick.
Most of the day was spent on beam development. Blue injection kicker
timing was adjusted, as was the collimation on the ramp. A study of the
collision component of Yellow Beam Decay was made. These studies, along
with a series of operator errors, resulted in a 1.5 hour beam development period
stretching into over eleven hours of unscheduled down time (from the
experiment's point of view). Booster C3 inflector and A3 RF station had
problems, and AGS RF K station tripped off and would not turn back on.
23 Jan 2008: The problem in the L-10 RF cavity was
fixed behind a store yesterday. A fill was lost yesterday when a
collimator (Yellow H2) lost its position and ran into the beam due to a limit
switch problem. After fixing the limit switch, the next fill was lost due
to beam losses at cogging. And the Yellow stochastic cooling kicker in
tank 3 failed to insert, so it was left disabled.
Good fill last night. APEX began at 5:00AM, except that the
experimenters (STAR) could not turn off their high voltage again.
APEX will continue until we start running again at 5:00PM.
22 Jan 2008: The vernier scan did not run yesterday until
19:00. Luminosities at STAR and PHENIX were higher after the scan than
before. The scan was performed on a store that had poor Yellow intensity
because of an error in filling.
Alex Zaltsman has requested an access to check on the L-10 cavity. He
believes that there may be a water leak in the piping, and is prepared to
replace the pipe if a leak is found. The repair will not take place
during this morning's access.
Both experiments reported that even though beam has been erratic this week,
they have recorded over 90% of the peak operating luminosity, so things are not
that bad.
21 Jan 2008: Gruesome weekend. The accelerator
is not behaving in a reproducible manner. There are many problems,
including instabilities that cause the emittance of the Yellow beam to blow up.
Emittance at injection to the Booster is about 5p
mm-mr, but by the time the beam reaches RHIC it is about 10p
mm-mr and growing. Under some circumstances the emittance growth is much
greater, only we are not clear about what those circumstances are. Power
supply problems in RHIC (bo11-qd3-ps) required an access to replace the faulty
supply. Controls had to restart two RHIC managers.
Angelika will do a vernier scan at the end of the morning store, around
11:00AM.
The AGS L-10 cavity tripped off again, as it has been giving problems all
weekend. When the L-10 drops out, the bunch intensity is low, and
sometimes "baby bunches" develop in deuterons.
In addition, Booster-to-AGS synchro is having problems during Au set-up.
Development of polarized protons has progressed well. Protons have been
brought out of the AGS. No polarization measurement yet in the AGS, but
that should be accomplished soon.
18 Jan 2008: PHENIX required an access after the
morning fill to replace and repair a HV mainframe for the MuID. Attempts
to repair the supply during an access yesterday were not successful.
During the access today, the Iris Scanner at PHENIX was not working, and Access
Controls had to reset the system.
There was a BLAM that shut down the beam from ~2:30 - 3:30 this afternoon.
17 Jan 2008: There was a major failure in the
Controls Database Server. Had to switch over to a back-up server while the
old server was rebuilt. Beam had to be off for the switch over.
While the database access was going on, both PHENIX and STAR made accesses.
PHENIX has a bad HV mainframe that they will try to repair.
There were first attempts to inject polarized protons from the LINAC into the
Booster without success.
BLIP came back on after fixing their sparking problems in Module #3.
Physics came back on after the Maintenance Day yesterday by ~20:30. Not
too shabby!.
16 Jan 2008: APEX ran from last night
until 8:00AM this morning. Start of APEX was delayed ~45 minutes because
PHENIX could not turn off their HV. Most experiments were completed.
Maintenance Day began at 8:00AM with the STAR repair of the ToF water leak
driving the schedule.
BLIP is still off due to sparking in LINAC Module #3.
15 Jan 2008:
Yesterday's attempts to raise chromaticity by one unit gave somewhat better
Yellow lifetimes and reduced coherence in the beams, but lower overall
luminosity. Then tunes at transition were increased slightly to compensate
for the tune shift observed across the bunches due apparently to electron cloud.
This gave somewhat more luminosity.
RF would like to put in the circuit for quad mode damping. It can be put
in behind the store, but it should be done at a time when no other changes are
made so that any effects of the damping can be easily determined. This
will be done for the new fill at 13:00 today.
The AGS Cold Snake is now cold. It has been hi-opt tested, and will get a
full power test tomorrow.
14 Jan 2008: Good running
over the weekend. Peak luminosities of 25 x 10^28 cm^-1 s^-1. have been
achieved.
There was a kink in the beam lifetimes which was due to the stochastic cooling
tripping off. It was fixed early Saturday, improving the beam lifetime.
Additional troubles with the stochastic cooling kicker drive. RF experts
had to be called over the weekend because the kicker for the Yellow longitudinal
stochastic cooling suffered a hardware failure. The same failure occurred
this morning between fills.
There was also a fix to the program used to compute the live orbit corrections.
This fix went in midday on Sunday.
There was a discussion comparing the amount of Au beam that makes it through
transition from the 2007 and 2008 runs. We are operating with ~10% fewer
bunches, and ~10% less intensity per bunch. Attempts to put more beam in
the machine lead to instabilities.
Sparking in LINAC Tank 3 has caused the BLIP intensity to be reduced by half.
They get turned off tonight at midnight to allow work on the Tank.
Hopefully it will be recovered during this week's Maintenance Day.
11 Jan 2008: Wolfram adjusted tunes to correct for
large losses in Yellow at the beginning of the store.
The emittance of the Blue Horizontal is very different from bunch to bunch.
Triggering on bucket 100 and integrating for ~50 buckets shows emittance of 5-10
pimm-mr, whereas triggering on bucket 150 gives 10-15. Yellow vertical and
horizontal shows emittances of 10-20 and higher. Where does the bucket to
bucket difference come from?
There was a long discussion of the reason for the high Yellow losses and short
Yellow lifetime. One possibility is electron clouds, and changing the
number of bunches would alter that. Turning down the voltage on the
storage cavity would reduce the momentum spread in the bunches. This study
will be done with the end of the morning store. The next store work will
be made on adjusting tunes to improve lifetime.
Todd has completed analysis of his Orbit Response Matrix (ORM) data, and is
ready to load the corrections. The purpose of the ORM is to find Gradient
errors. Todd wants to load the correction. Thomas does not want to
load them on Friday. The compromise is to load the new settings next week
during the APEX when Mai Bei will take an orbit measurement before and after the
load to quantify the improvement.
It is Friday, so we are looking for the Blue Noise. Kevin is ready to look
at it using the LFShottky. Paul has contacted LIPA you are looking at
noise on input power. Todd has ideas for isolating the source of the
noise.
10 Jan 2008: APEX running continued until 18:30 last
night, thanks to the allowances from both experiments. Half the APEX
program was completed. The remainder of yesterday's schedule will be added
to next week's APEX.
Immediately after APEX, there was an access last night to allow PHENIX to
remove their converter from the beampipe. During the access, Mike Brennan
made adjustments to the stochastic cooling. The power supply for yo4 also
needed to be replaced. STAR made use of the access as well.
A good physics store followed the access, starting at ~20:30. When
filling the next store, beam losses dumped the beam, followed by a quench.
While trying to recover from the quench, the QH8 magnet power supply in BtA
failed at 4:55. As of 9:00AM we are still trying to bring QH8 back on.
While under LOTO for the QH8 work, the polarimeter electronics at C15 was
powered on. The RF station C amplifier was also replaced during this LOTO.
9 Jan 2008: APEX is underway. Problems with the
C7 Booster bump delayed beam until ~9:00AM.
We have not been able to repeat the good luminosity fill of yesterday morning.
At least one of the three intervening fills had an instability in Yellow.
Mike Brennan needs to work on the Landau cavity in Yellow. He also wants
to adjust the voltage on the stochastic cooling to address the problem of
satellite bunches. Carl Schultheiss needs time to work on the transfer
function of the dipole. Mike Brennan needs to make an access to the 12
o'clock hall to check on the alignment of the stochastic cooling plates.
This will take 5 minutes. If the plates are out of alignemnt, then another
half hour access will be required to bring them into alignment.
PHENIX has been running since yesterday with a brass sheet around their beampipe
to calibrate the photon conversion rate. They need to take that sheet out
now, requiring the East Carriage to be pulled back. This necessitates a
half hour access at the end of APEX, or at any time there is a failure during
APEX. Once they get back on, and get a store going, they will take a 1
hour long magnet-OFF run to measure the location of the East Carriage since it
will have been moved. This will be accomplished during the last hour of a
store, perhaps the first store after APEX. It should not impact either
STAR or RHIC.
STAR will take the opportunity to make an access during the PHENIX access.
8 Jan 2008: Instabilities on the ramps required
beam intensities in both Blue and Yellow to be reduced throughout the night
until finally a good store was achieved at ~9:00PM last night. Then at
4:00AM a store went up with moderately good beam intensities and great
luminosity.
STAR needs a 30 minute access to replace a VME power supply. PHENIX will
make use of the access to install a converter. This will occur at the end
of the morning store at 10:00AM.
7 Jan 2008: A problem with the agsrhichome server
has disabled outside access to the rhic web sites. This server has been
powered down and ITD is looking into the problem.
Late Friday we moved to 93 bunches. This led to higher luminosity, but
tune problems have kept us from realizing these gains.
A series of problems has made for rough running this past weekend. The
noise source that causes blow-up in the Blue emittance returned on Friday night
and then switched off again Saturday.
Tune changes from above the diagonal to below it have reverted to the earlier
tune. Currently we do not have the Blue and Yellow tunes on opposite sides of
the diagonal.
Scheduling issues:
4 Jan 2008: Gold injection has been a problem since
the flip of the PHENIX magnet polarity.
Power supply problems with b02-qgt required a new jump card to fix Blue losses
at transition. An access was made to Alcove 3 to replace the card.
Last night's fills were interrupted by two RF problems. Blue Storage
cavity 3 (BS3) had trouble turning on because of low temperature interlock.
Yellow Accelerating Cavity 1 (YA1) had a bad relay in the PASS indicator.
An access was required to turn on the Ignitron of BS3. When they were
brought back on, beam could be filled and stored.
3 Jan 2008: Getting going again after the maintenance day was not smooth. Retuning
of both beams was necessitated by the polarity flip of the PHENIX magnets.
Pressure in the PHENIX straight section bi8-cc-pw3.1 increases by nearly 3
orders of magnitude. The PHENIX IR vacuum shows no pressure bump.
After getting a 6x6 bunch store in around 10:00PM last night they attempted an
87x87 bunch store. Yellow efficiency is down, and Blue exhibits a drop
when Yellow is injected.
STAR was unable to work on their water leak because of hydraulic problems in
the pole tip extraction.
PHENIX did the magnetic field flip, but were unable to fix the Pad Chamber
DAQ problem.
Baby bunches in AGS. When Peggy is able to resolve this problem, the
experiments request that the beam be dumped and refilled.
Booster Main Magnet Power Supply (MMPS) problem is keeping us from refilling
in the morning.
2 Jan 2008: APEX worked through the night.
Maintenance Day began at 8:00AM.
Change in the tunes improved the efficiency of both the Blue and Yellow
transmission.
Blue Accelerating Cavity #1 went down on Tuesday. A driver in the ring had
to be replaced. Beam was lost setting a BLAM alarm.
There was no work done to study the correlation between Blue vertical emittance
growth and Air Conditioner cycling. This will be attempted later this
week, probably Thursday.
STAR has the FMS trigger implemented as of Monday. It is still being
developed. They plan to complete the study of it today, and it will be
added to the list of Physics Triggers.
PHENIX has collected more than half of the total luminosity they require for
d-Au running. They have reversed the polarity of their central magnet.
This will require retuning of the orbit, so a 6 bunch ramp will be brought up
first to make these corrections. In the past this required a change in the
orbit and coupling. This year feedback on the ramp complicates matters.
Last year we never recovered the same rates as before the field reversal.
There is also a question of whether or not the electron cloud will create vacuum
problems in PHENIX. we have not seen bad vacuum effects in PHENIX yet this
year.
31 Dec 2007: Good running over the weekend. The
running sum of luminosity is now corrected for accidentals.
Saturday morning had fills with very bad Blue lifetime, due to a noise source
discovered ten days ago when looking at the LF Shottky signal. It appears
as if something is turning on and off, but what is it? Thomas, Wolfram and
Peter looked but were unable to find a source. There is no intensity
dependence. IPMs show the same emittance growth. Studies of the slow
logs and pscompare showed no guilty party.
At the end of last week, we started doing automatic orbit corrections every
half hour during physics stores. This improves the beam lifetime and
reduces the experimental background each time.
Christoph changed the chromaticity to increase transmission efficiency.
PHENIX reports the North side backgrounds are getting worse. They
request a new collimation or tuning effort. This is new because they have
traditionally been troubled by the Blue beam hitting the South Muon ID counters.
PHENIX will want to reverse their magnet field on Wednesday since they will have
accumulated half of their total luminosity for the run. They also have a
Pad Chamber problem with the electronics for PC1. This will be addressed
on the Wednesday Maintenance Day.
STAR saw a low pressure front that set their TPC into Purge mode since it
couldn't keep up with the dropping barometric pressure. They are still
struggling with the high level triggers for the FMS. Other than that they
have been accumulating good luminosity. STAR plans to open their magnet on
Wednesday. It has a small water leak. Fixing this will be an 8 hour
job on the Wednesday Maintenance day.
Access Controls will remove the EBIS chipmunk from the LINAC UPS, decoupling
AGS running EBIS.
Polarimeter work was done on the Blue electronics, cycling the power to the
Blue and Yellow HV supplies and the DVM. This fixed the communication
problems we were having last week. Yellow preamps/shapers will be
installed during Wednesday's Maintenance Day. Currently there is a
calibration run underway to understand the three different electrostatic
shielding designs being tested for the preamp housing. AGS cold
snake is being cooled down already. LINAC sections are in preparation for
the polarized proton running. LINAC will be ready in ~two weeks. AGS
polarization work will need to start ~two weeks ahead of pp running.
APEX work begins at 8:00PM tomorrow night, continuing to 8:00AM? Wednesday
morning when Maintenance Day begins.
28 Dec 2007: After running good physics stores all night,
STAR needed an access when their magnet tripped off..
Vacuum group flashed the TSPs in the Blue 7 o'clock warm straight section at
4:30 yesterday afternoon. They report that there was no observable
improvement in the vacuum. PHENIX reports that there was no noticeable
improvement in the background rate. So there is no plan to implement a
flashing routine.
The store that came up at 17:00 yesterday had high backgrounds from Yellow at
PHENIX. MCR had to pull out the collimators (primarily the Yellow vertical
collimators) a bit to optimize running.
Stochastic cooling developed a problem It appears it is
operating at the wrong gain, making for slow cooling. Mike Brennan found a
bad fiber optic connection and "fixed" the problem at about 10:50AM yesterday.
27 Dec 2007: The work done yesterday to reduce the
beta* in the Blue beam went very well. The luminosity has been increased
without significantly increasing the backgrounds, at least according to STAR.
PHENIX has still to report. Yellow lifetime appears to have deteriorated
somewhat. The beam intensities per bunch in the AtR are 1.2E11(Blue) and
1.0E9(Yellow). Previous running had seen 1.4 in both, so there is some
room for improvement.
PHENIX recovered the North Muon Arm during an access late Wednesday.
Mike Brennan reports that the stochastic cooling amplifiers are maxed out,
showing that there is something wrong and maybe the cooling may not be
functioning at all.
Vladimir points out we have not done a 93 bunch store or made use of the fancy
gap store yet. He also notes that the Yellow and Blue tunes drift towards
each other, This causes emittance blow up. Usually it is the Yellow
tune that drifts.
Christoph will incrementally try to improve the efficiency at transition by
playing with the chromaticity.
Mike Mapes has flashed the TSPs in the warm section between 8 and 9 oclock in
the Blue ring. After the next fill (`16:30) he will flash the TSPs between
7 and 8 oclock, just upstream of the PHENIX Muon ID. If this makes a
measurable change in the background levels, then flashing of these TSPs will
become part of the routine.
PHENIX and STAR have requested a brief 15 minute access after the 16:30 beam is
dumped. This will coincide with the vacuum work.
26 Dec 2007: There was a high-level RF problem this
morning at the beginning of the APEX. The Yellow Accelerating cavity, YA1
lost voltage. This follows yesterday's continuing RF troubles with
the X1, X2 YS3 cavities. When the X2 common cavity drops out, some of the
Yellow beam and much of the Blue beam is debunched. Several times this
occurred. There was a beam loss that tripped the BLAM early this morning
allowing APEX to start early (5:30). It also caused a real quench.
Also, the continuing problem with the power supply for the Sextupole yi3
reoccurred. It lost its fiber optic control card for its power supply.
Replacing the card returned functionality.
APEX will study the effect of squeezing the beta* in the Blue beam today.
PHENIX reports a failure in the North Muon Arm. A 15 minute access
resulted in smoking a transistor in the crate supplying low voltage to the North
Muon Arm. A longer access has been requested, >1 hour, which will take
place at the end of APEX or during any extended downtime during APEX.
Peter Ingrassia reports that the integrated luminosity for Christmas Day exceeds
that of any day recorded so far.
24 Dec 2007: Trouble with the RHIC RF. BA1
(Blue acceleration cavity #1) lost a neutralization loop in the power amplifier.
It needs to be re-soldered back on. It is fixed and we are back running.
Sextupole trip caused loss of most of the Yellow beam. After running
without the Sextupole for an hour, the beam was dumped after polling the
experiments.
There is a bad gate interlock (at the switchyard upstream escape gate) that
needs tensioning. This will be done behind the first store this morning.
Blue emittance growth problem solved by unplugging a white noise generator in
the RF system.
Lifetime at injection is bad and needs to be fixed. Lifetime at store is
fluctuating, apparently due to variations in power supplies or orbit.
21 Dec 2007: Discussion of the reasons for all the 6
lost fills yesterday. The first two were due to ripple on the power
supplies. Carl pulled a card to fix after the first fill was lost, but did
not complete the repairs until after the second fill was lost. The third
and fourth fill was lost when a procedural error resulted in a sequencer being
run with no separation bumps. The fifth and sixth fills were lost due to
an instability produced when the Landau cavity voltage comes on.
Yellow efficiency remains good, but Blue efficiency drops throughout the ramp
and transition. It is not clear why. The same losses appear in 6
bunch ramps, so it does not depend on the number of bunches.
There is a smoke alarm in the Blue QLI system. Yugang would like to
reprogram the alarm so that it doesn't take down the beams. There was a
question of whether the alarm time could be lengthened, since it appears to be a
transient causing the false alarm, and Ray Karol needs to make this assessment.
The repair would keep beam off for approx 2 hours, so it was put on hold
awaiting any unscheduled downtime. Paul will find out whether the
reprogramming can be accomplished over the weekend or holiday, should the
downtime occur then.
There has been a lack of communication between the RF group and the OC
regarding the status of the RF. The RF has been left in CW mode causing
unnecessary delays in getting beam going. This communication needs to be
improved, and perhaps have an alarm in the MCR showing the RF status.
STAR needs to replace a bad HV supply card. They are waiting for delivery
of the new card. As soon as the new card arrives, then at the end of the
subsequent fill they will make a 15 minute access to replace HV cards.
Mike Brennan needs a half hour between stores to work on RF phasing. This
work can probably occur during the STAR access.
The OC worked with PHENIX this morning to study correlations between PHENIX
background scintillator S2 and the PHENIX Muon ID currents. There seems to
be fair correlation, so the S2 rate will be used to tune up the PHENIX
backgrounds.
20 Dec 2007: Recovered from Maintenance Day with
first store at 9:00PM. Began with a 6x6 bunch store, with an 87x87 bunch
store for physics at 10:40PM.
Using the new dAu81 lattice with beta* of 0.8 in Yellow. The new lattice
appears to have a smaller emittance in Yellow than Blue at the beginning of the
store.
STAR lost their BBC counters and needed to make an access (circa 5:00AM) to swap
a card, so they do not have an estimate of the background rates with the new
lattice. Vacuum near STAR is improved somewhat, down to 1E-10, down from
2E-10. PHENIX reports that the backgrounds look somewhat better.
Travis tuning the Blue collimators had to pull them out from their set-points to
reduce the PHENIX backgrounds reported by scintillator S3.
19 Dec 2007: Maintenance Day is underway..
Using dAu81 lattice (reduced beta* in Yellow), obtain 85kHz of ZDCs for per
bunch intensities of 0.85 for both Au and d. The measured beta* in Blue
was 0.91/0.93 H/V, and for Yellow, the beta* is 0.81 for Horizontal.
Mei did these studies. The displacement of the beta* minimum from the real
interaction point is 0.8m in Blue and 0.3m in Yellow. This could be a big
effect, since a 1 meter displacement has the effect of reducing the effective
luminosity by a factor of 2.
Experiment background rates were measured as comparable to the larger beta*
lattice.
There is a debate about how to best scrub the vacuum out.
18 Dec 2007: Ran stores last night, but lost the Blue
clock causing PHENIX to be unable to take data. Also lost the ability to
dump the beam cleanly. So when the beam was dumped, there was a real
quench as well as a BLAM 90% alarm (90% of hourly radiation limit reached,
shutting us off for an hour). It appears the blue cogging was lost when
the RF clock synch stopped.
There was an attempt to scrub the vacuum yesterday afternoon by repeatedly
filling at highest intensity at injection energy. Monitoring vacuum levels
at STAR showed that scrubbing did not succeed in raising the vacuum pressure
even as much as standard colliding beam. Further "scrubbing" will be
accomplished by production stores.
PHENIX lost a Low Voltage Power Supply when they lost the beam synch.
They needed to make an access to replace that PS. STAR also took advantage
of the BLAM down time to take a short access.
STAR is still concerned about the background rates. Their minimum bias
trigger is only about 40% efficienct (60% of these triggers are accidental).
As soon as the morning meeting ends, dump the current store, then Don Bruno
will load the new lattice. While digesting the results of that work, there
will be a new store put up with the reduced beta* in Yellow. Then there
will physics stores until the APEX starts at about 8:00PM tonight (or later.
All of the work Don does this morning will be subtracted from the front end of
the APEX schedule).
17 Dec 2007: New operating instructions for injecting gold
beam. Now we must just kick out the "pilot" bunch, leave three more
bunches in booster. We must always have extraction going so beam is dumped
in the W-dump at all times, Dumping beam even at low energies is a
problem.
One difference between last year and this year is that last year we did not
make use of early turn off. We also forgot to turn on the dump bump a few
times, bringing the beam to a bad location outside of the AGS beam dump.
Have changes in collimator settings improved things? The collimator
work brought down several stores on Friday.
Instability re-appeared late Sunday night. Increasing the octupole
strength helped previously, but has deteriorated since.
Rebucketing in Blue has deteriorated also.
Gap cleaning kicker in Yellow is ringing and effecting bunches outside the
gap. Angelika thinks that with stochastic cooling on, we don't need gap
cleaning. She suggests turning it off.
Yellow abort kicker prefire brought down a store on Saturday at 12:20. This
coincided with a power glitch the blew fuses in PASS, blew out an ODH sensor,
tripped several radiation sensors.
STAR sees ZDC singles rates of 2MHz. ~60% of this is not collision
related. STAR cannot run like this. Will need to spend some time
today immediately after the morning store ends(11:00AM) to scrub the vacuum.
A crow bar fault on b4 has brought down several stores. Don wants to
hang a 16 channel data recorder on the power supply to try to catch the problem
in action. May need to swap in a replacement PS.
There is a discrepancy between the emittance measurement that vernier scans a
IPM gives. This discrepancy has existed for years. The
emittance measured in the ATR is a factor of two larger than the results from
either the AGS or RHIC IPMs.
PHENIX reports that they will need to make a brief access on Wednesday to fix
communication with the MU Tracker. PHENIX also has a broken railing on their wheelchair ramp. It is a
safety hazard.
15 Dec 2007: Another vacuum leak in the AGS. This time in sector I17. It is the same symptoms as the previous two leaks; i.e. burned kapton. The leak occurred after 10 minutes of Gold tuning with "Early Turn off" turned On. The leak was repaired with spray sealant.
14 Dec 2007: STAR will begin the day with a two hour 6
bunch store.
After the 6 bunch store, PHENIX wants to do a vernier scan with Angelika.
The stochastic cooling is causing the central bucket to diffuse into the two
neighboring buckets. On the first production store Mike Blaskiewitz will
look into tuning up the SC.
We are currently at 87 bunches. Per bunch intensity is at 1.1E9 for Au,
and we know we can grow to 1.2 or even 1.4. The deuteron intensity is
close to 1E11 and also has room to grow. The easiest way to grow is
probably to add bunches. The limiting issue is likely to be the growth in
background rates. The residual gas pressure is up, and the background
rates scale linearly with pressure. Will try 1.2E9 in Au today.
The STAR 6 bunch fill took longer to get started, and took longer to accumulate
enough data to be completed early. The 6 bunch fill did not dump until
3:00PM. This meant that the PHENIX background study did not get started
while Angelika was available. We will have to revisit the PHENIX
backgrounnd study on Monday 17 Dec 2007. Difficulty filling slowed the
start of the physics fill at 16:00.
13 Dec 2007: During the APEX work yesterday, there
was a reduced beta* lattice developed, but not completely evaluated.
During machine development today, we will try to finish the evaluation, and try
using the reduced beta* lattice in Yellow. If this works well, we will try
to reduce beta* in Blue as well.
PHENIX wants to coordinate with Angelika on a vernier scan to understand the
backgrounds.
Earlier this week, Todd did orbit corrections for the quads. Do these
corrections scale with the new beta* lattice? This needs to be tried.
Mei will do orbit and dispersion measurements at store.
Beta* is measured horizontally in IP8 and vertically in IP6. The BPMs were
not available for the other measurements, This needs to be completed
today.
Friday morning, STAR would like to do a 2 hour study with 6x6 bunch store.
12 Dec 2007: APEX work all day.
STAR requested an access for 10 minutes to cycle the power on a crate before
going back to physics at the end of APEX running.
There has been a discussion of how to improve experimenter's access.
PHENIX did an access last night and after half an hour of trying to get a key
release, they had to have MCR provide a gate watch. STAR access was
extended because of an insufficiency of keys. There is a plan to shift the
3 pp2pp keys over to STAR usage at least until the beginning of pp running.
11 Dec 2007: Increased the bunch intensity from
0.6E11 to 1E11 deuterons in Blue. Vadim noticed that the vertical tunes in
Blue and Yellow were too close to each other. Now they are more separated,
allowing for the increase in Blue. And the number of bunches has increased
to 87. Now ZDC rates peak at ~90kHz.
Todd has a study he would like to do with the last hour of the morning store.
SteveT is ready with the new small beta* lattice. Don will attempt a
couple of ramps today at the end of the morning store before refilling.
Behind the ramps, Mike Brennan wants to do some RF work. Problems with a
sticky tuner. Blue Storage problems also.
10 Dec 2007: The Au beam has been dumped inside the AGS
whenever we have been tuning. There is evidence from the loss monitors
showing that the beam has been hitting the J-17 section, just where the vacuum
leak occurred. The practice known as "early turn-off" has been to stop
accelerating the beam during tuning, but the magnets continue to ramp, so the
beam spirals into the inside of the pipe. The same practice is true when
bad bunches are dumped internally during RHIC injection. Is the berm
capped over J-17? It has soil-crete over top of it.
SteveT needs to try reducing the beta* from 1 (or 1.2) to 0.8 or even 0.7 today.
DonB needs to try a the new ramp without any beam in the machine before
repeating it with beam.
08 Dec 2007: Vacuum leak in J-17 straight section. Unable to fix it with Glyptol, too big a leak. The kapton is burned through, just like the L05 section. Pulling out quadrupole to replace vacuum section. Expect to get back on line by ~8PM. The bigger problem is "Why are we developing beam-related leaks in the pipe?" This is going to require a careful study of loss monitors, and understanding of how this could be happening.
07 Dec 2007: Two RF cavities tripped off last night, so
RHIC filled with just a single cavity. Then that cavity tripped off,
leaving just the common cavities. It was a very good fill. The RF
Igniters (Ignitrons) must be warm enough to fire, and the AC in the RF buildings
is too cold.
Nominal per bunch intensity for good running is 0.6E11 d per bunch and 1E9 Au
per bunch
STAR report high background rates, which seems to correlate with vacuum
levels. PHENIX also sees high background rates, although primarily in the
Blue side.
Do steering studies during the day today, but otherwise just run good stores
all day and weekend. Experiments would not take data during the study, but
it would not effect the store, and physics running could continue before and
after. Doing a vernier scan will allow both experiments to check how much
of the background rates are due to collisions.
Chromaticity was measured up the ramp. Two sets of measurements of
tune, with the radial offset on and off, to give the chromaticity.
PHENIX wants to do a zero-field run at the end of this fill. It was agreed
to by RHIC and STAR.
Mike Brennan wants to bring on the Landau cavities for the next fill. It
was agreed to.
06 Dec 2007: Maintenance day Problems during the maintenance day continue to plague us.
The beginning of the maintenance went well, with Booster and AGS back on before
noon, and PS work in testing mode, as requested. But the work on the
breaker at Building 929 was initially delayed because the parts had not be
delivered in time. They finally arrived at BNL at 10:30, coming to C-AD
after lunch. After some discussion, it was decided that since the breaker
had been performing in a stable manner for the two week period during which
measurements had been taken, so we could postpone the work for two weeks, until
the next maintenance day. Two hours later Jon Sanberg reported that the
breaker in Building 929 had increased in temperature by 11 degrees in one day.
At this point the risk of postponing the work seemed too great, and at 2:30 PM
the work on replacing the breaker began. The actual breaker exchange took
only about 15 minutes, but with all the end effects, we were ready to be back on
by 18:00. We began with a 6x6 store which went well, then attempted a
78x78 bunch store. There was a QLI followed by another QLI. Then
MMPS tripped off, and RF tripped off, followed by a couple of Permit Interlocks
being pulled by Loss Monitors. It took until 6:00 AM the next morning, ~12
hours late, when we managed to put up a store following the Maintenance Day.
We will have to think hard, basing our decision on data taken from all the
maintenance day, day after and week after maintenance days to judge whether
Maintenance Days are beneficial or not,
The beam needed a chromaticity jump to +2 at
injection to get good efficiency, but then -2 when the beams are brought into
collision. Clearly something is wrong. We need a knob to adjust the
chromaticity.
The Yellow Landau cavities are not yet on. Mike Brennan thinks that at
these intensities, we would benefit from turning them on.
Without collisions, the beams are at an unstable chromaticity. Bringing
the beams into collision then causes poor lifetime and large loss rates.
05 Dec 2007: Get ready to run 78x78 with reduced deuteron
intensity. The TSPs were activated yesterday during a ramp, so it is
important to see if this brought down the background rates in STAR.
One of the studies last night was to prepare 3 fills with d intensity of 1.0,
0.8 and 0.6 x 10^11 ions. (Keith was able to inject 1.4E11 for one store)
But ya1 had tripped off. Ron Schroeder was providing a noisy BBQ signal to
increase the Blue emittance, from 14-15 to 17-18. Comparisons of the
lifetime of the three different intensity bunches show that over a twenty minute
store, there was no difference between the three different bunch sets.
The BtA foil drive is not functioning well. The current foil cannot be
moved. We ran last night with the Au foil left in for D running.
Keith reports that he can't tell the difference in the D beam. It may give
a slightly increased emittance in Blue, which is the direction we want to move
anyway. It seems as if running until the next maintenance day (2 weeks)
without fixing it is good. The drive is binding, but can still be
moved by hand. Continuing to move it will cause it to seize. .
04 Dec 2007: If we are running polarized protons this year,
then we need to pay for the LINAC startup and operations reimbursing BLIP, but
this is not covered as part of the continuing resolution funding. So we
need to come up with a plan for how pp running will be decided in early January.
There was no discoloration or burning in the stainless steel section of L05
vacuum section that was replaced last week, only in the kapton itself.
Vacuum is ready to start activating the TPSs once or twice a week during
downramps. Vacuum wants notification if scrubbing begins in
order to change the setpoints at which the protection valves shut automatically.
They setpoint is currently set at 10-7 although vacuum rises have
been only 10-10 or so last night.
Leif will check the kapton on the new L05 vacuum section to see if there is any
new damage.
PHENIX wants to turn off their magnet at the end of a store (it may have
occurred this morning) to measure straight tracks. STAR may be asked to
turn off their magnet at the end of a store to allow the electron cloud to
disperse without raising the background levels to unacceptably high levels.
03 Dec 2007: AGS Klixon tripped off the MMPS.
Probably a controls issue.
Stochastic cooling in Yellow doesn't seem to have any effect on beam lifetime.
It appears to be reducing the bunch length, but not improving things. It
is as if closing the aperture by putting in the jaws is reducing the momentum
spread, but the cooling is not taking place. It also appears as if Blue
bunch length is growing with and without cooling on in Yellow, but Yellow is not
growing even when cooling is off. Quite a puzzle. It is possible
that BBQ is not off, and so could be injecting noise into the system, heating up
the beams. This is not a tune space issue. It is a real transverse
emittance growth colliding with betatron aperture. We need to get the IPM
working. The problem is not small aperture, but rather large beam
emittance.
Yun and Vincent applied chromaticity corrections yesterday, but it made beam
losses worse. They were hampered by ramping speed. The chromaticity
corrections reduced the chromaticity beautifully, by a factor of two or more.
But this did not improve beam lifetime. So the chromaticity corrections
are working well, but that is not the problem with the beam.
Use the Shottkey to measure the beam emittance. So begin by doing
the 18 Amp measurement with Q2, then Todd wants 4 hours to do Orbit Response
Matrix (ORM) measurements, turn on the logging of Shottky data, verify that the
BBQ is really off, verify that gap cleaning is off. Try running with no
rebucketing, no gap cleaning, no stochastic cooling, no collimator, and look at
the beam lifetime first without bringing the beams into collision. Then
start collisions. Then slowly try turning on the apps, with cooling last,
to see which app is responsible for short beam lifetime, and large beam losses.
PHENIX started taking data on Thursday morning at 00:01 AM. STAR began
data taking on Saturday morning 8:00 AM with minimum bias triggers. PHENIX
sees large backgrounds coming from the tunnel on the Blue beam side, as shown by
the MUID wire currents
There was a suggestion that the reason for the poor lifetime is colliding beams
with different emittances. It was an effect noticed first in pp, when one
of the beams had a large emittance, and the other had small, then the big beam
tended to blow up even faster.
02 Dec 2007: Stochastic cooling was on for the stores
last night, but it appears that it doesn't have much of an effect on the
lifetime. Both experiments complain that when the cooling gets turned on,
it produces a large burst of noise. It appears that the Yellow lifetime is
governed by something other than beam beam interactions. Todd is concerned
about the vertical chromaticity in Yellow. There needs to be a
chromaticity measurement in Yellow. Vincent doesn't want to do the
measurement with 59 bunches in the machine. Try to do a 6x6 store for the
measurement, if we can get the right people in. If we have a large
non-linear chromaticity in Yellow, that may prevent us from fixing the linear
part. Thomas wants to make sure that the BBQ is on. It is suspect,
and so should be shut off until it can be checked. What does the vertex
distribution look like? It should make a big difference in how large the
interaction vertex is. The most recent store looks like it had several
kinks in the vertexing efficiency, as if the cooling either came on or went off
during the fill. The chromaticity needs to be very small in order for the
cooling to work without large losses. The machine is unstable with
negative chromaticity causing emittance growth. There was a store with
good Yellow lifetime on 30 Nov, so try reloading the stones from that store and
see if the store repeats.
There is currently a 59 bunch store being set up, so we will call in Vadim, and
when he is in, we will dump the store and refill with 6 bunches.
27 Nov 2007: Since we declared physics yesterday at 400PM,
there will be machine development only on the day shifts 8:00-4:00 weekdays.
No weekend machine development. APEX will start next week, 5 December 2007.
10, 12 and 2OClock Iris scanners will be disabled. Will install 4OClock. Need to check with Paul to make
sure it is on the maintenance schedule.
Vacuum is getting ready to replace the AGS L05 section of vacuum pipe.
MikeB was work to do on the pickup coils in the 12OClock and 1OClock halls.
New tubes for the 200MHz cavity are due in 17 Dec 2007.
EBIS field trip to Germany tomorrow to learn about the solenoid failure.
Pulsed Power wants to solder the connections to the ??? And install the Yellow rechagrge circuit.
Reconnect the L20 output cable. F3 PLC modifications Abort Kicker oil leaking capacitor will be replaced.
Instrumentation sees improvement in the IPMs, so will retrofit the others during the run, one at a time.
John Morris needs to do work on getting the replay working.
It is possible that the L05 vacuum chamber was misaligned, and that was the reason for the beam
hitting it causing the vacuum leak last week.
Several 56x56 stores were lost last night when there was a Beam Permit
pulled followed several seconds later by a Quench Link Interlock, not a quench.
Attempts to inject with beam at intensities of 1.06E11/1.09E9, 0.98E11/0.63E9,
0.86E11/0.66E9 in B/Y.
Wolfram and Peggy will begin by working on aligning the abort gaps in Blue
and Yellow.
MikeB found that at the frequency jump, two of the storage cavities suffer a
drop out in voltage. May request a high intensity ramp to see if it gets
worse, Also would like to try the stochastic cooling in Yellow before tomorrow's
maintenance day. Want to insert the three tanks and scan the beam to see if
the beam is centered in the BPMs. And stochastic cooling is ready to go.
Can be accomplished with a 6x6 ramp. Need about an hour, immediately
following Wolfram and Peggy. Assuming it works well, it could be turned on tonight.
Check aperture with 6x6 ramp.
If BBQ is working, both Blue and Yellow tunes should be flat. Currently
only the Blue is flat. Why can't we make Yellow flat? Coupling data
is good enough to do a feedback ramp. Need about 3 hours to do this.
Mei wants to measure the beta* at store in Yellow, since yesterday's
measurement was compromised by the rebucketing. Will take about 20 minutes
at injection. Mei also wants to do orbit measurements in Blue at
injection.
Al wants to get the Yellow feedback ramp and replay going.
Do orbit, coupling and chromaticity corrections.
Wolfram wants to measure the tune shift from the front to the back of the
bunch trains. .
Wolfram thinks it may be important to do some scrubbing. Vacuum wants to be notified the day
before, and requires 3 hours set-up. Possibly do scrubbing on Thursday.
26 Nov 2007: Tune injectors up to 1E11 for deuterons and 1E9 for Gold. Intensity had
been dropping throughout the weekend. It appears to be improved now.
Tune jump in Blue at the start of ramp whenever the replay file is used.
There are three strategies for dealing with this, but all involve removing rapid
tune changes at the beginning of the ramp, either by regenerating the ramp file
without tune corrections, filtering out the early corrections, or modifying the
software to apply the corrections slowly.
Rebucketing in Blue is required, giving to big longitudinal emittance.
MikeB would like to trouble-shoot the RF to see if that is the source of
the rebucketing problem. The rebucketing display also fails to show both
beams on occasion.
There appears to be coupling near the end of ramp in Blue. Yun will
look into it today.
Radial excursion around transition in yellow has gotten large again.
MikeB noticed that Blue Orbit appears to show a bump when the Yellow is going
through transition. Blue goes through transition 8 seconds before Yellow,
so Angelika thinks the display is correct, and the excursion in Blue is in time
with passing through Blue transition and not related to Yellow transition.
It was agreed that today beginning with the evening shift, the experiments
will have collisions from 4:00PM through to 8:00AM, with the day shift going to
beam development.
First try the new replay file, then fix rebucketing.
Last night we attempted more bunches, but had to retreat because of losses at
t260.
Collimator controllers failed over the weekend, causing a beam loss when a
controller walked a collimator into the beam over a one and a half minute
period. But the current beam excursion of 1.5mm is too large for
collimator tracking.
PHENIX has a safety issue with their front door. Fred Kobichuk has
called in the work order, but nothing has been done yet. The entire door
frame will probably have to be replaced. PHENIX wants more beam intensity.
Total beam is down ~2, and vertex efficiency is down a factor 4-5. Both
need fixing.
During a fire alarm in Bldg 1010B (apparently caused by a faulty smoke alarm)
Steve Jao made an access to TTB to replace a chip in the Current Transformer
amplifier in Section 29. Earlier attempts to fix this beam current monitor
did not respond to any of the easy fixes.
In response to a request from the vacuum group, scrubbing will not be done
until it is required. And currently it is not required, so there is no
plan for scrubbing.
Alex wants us to reduce the gap volts in the Band 2 system. This
applies to both deuterons and gold.
Maintenance day planning: PHENIX has a
light day's work on power supplies. PHENIX also has a non-functioning
pulser in the South Tunnel that needs repair by Steve Boose. Also some
routine Air Conditioning maintenance that Charlie Pearson is doing.
Charlies' folks are requested to check in with PHENIX Control Room before making
access and when done. STAR says they will use the entire 8 hour access for
unspecified work. There is an access request for the 12OClock Hall to
replace a pump on the JET.
We have 56x56 with bunch intensities of 1E11 Blue and 0.6E9 in Yellow with
good rebucketing (the best so far). Landau cavities have been retimed.
Tonight we will
fit in three 7 hour stores between 3:00PM and 9:30AM tomorrow morning.
APEX begins on Wednesday 5 Dec 2007. Since the plan is to do a
maintenance day the same day, APEX will start at 00:01 AM.
25 Nov 2007: Beam intensity has been dropping throughout the weekend, for both Bleu and
Yellow. Need to investigate loss in intensity.
Plan to work on injecting more bunches. Currently running 68x68 but
there are store files for 78, 87, 93, 105 bunches so it should not be a problem
to test them out.
Work on vernier scans in both STAR and PHENIX will be done. The old
operating point was very close to the peak for both.
PHENIX vertexing efficiency has been low; ~37% instead of the expected 50%.
STAR sees the interaction diamond with an RMS of ~70cm.
24 Nov 2007: Blue Landau cavities were turned on, curing
the dancing bunch problem.
Communication problems with a flakey chipmunk that had to be replaced, and a
cryo temperature reading at 4OClock.
Ran a 6x6 fill for a STAR vertex check. Vertex is consistent with being
good, but RMS of vertex is so broad (70cm) that it is impossible to determine
the vertex with any precision.
There is a tune jump at the start of the Blue ramp when in replay mode.
This will need to be addressed.
23 Nov 2007: PHENIX took a magnet-off run. STAR
took a run with their TPC HV on.
The Yellow V1 collimator motor was fixed after it was found to be wandering off
during fills.
22 Nov 2007: Rebucketing problems with both rings.
Beam is "dancing" on the ramps. Plan to turn on the Landau cavities to
help with this.
The is coupling between the Blue and Yellow beams at injection, such that there
are beam losses in Blue when injecting into Yellow.
Tried six ramps with 37,56, 68, and 87 bunches.
21 Nov 2007: The experimental access began today a
little early (8:30AM) Both are at restricted access. In
addition, there will be an attempt to fix a broken IPM, and a swap of a power
supply card at yi10-tq4.
Q89 in yellow is exceeding the current limit of 275Amos. It is at 289A.
Last year this supply gave a lot of problems at this current. Lowering the
limit to 275A made most of the problems disappear. The hall is cooler this
year, so it may be OK now, but Don would prefer to operate at lower current.
Instead of operating at such a high current, we need to develop a new lattice
that does not require such a high current in q89. SteveT needs to work on
this during the access today.
The rebucketing done earlier in the week appears to need to be redone.
This was a "data management" error. Now that we have rebucketed beam, it
is time to turn on the stochastic cooling. There was a lot of spillover,
gold more than deuterons. There is work to be done to get tighter bunches.
Background rates were about 4 times as high as they need to be.
So it is clear that we need to recheck the collimation, especially if we redo
the lattice.
Gap cleaning was working yesterday, but was not turned on in this morning's
stores.
20 Nov 2007: Three separate ring accesses yesterday. Blue
abort kicker had problems related to the smoke detector. PHENIX had work
on its Iris Scanner. STAR did trigger cabling work.
yo8-tv5 (IR8 separation bump) failed and had to be replaced.
Intermittent Quench Ground Current warnings
Spent a couple of hours trying to decouple yellow. First attempts did not
work,
PHENIX is having trouble with the Blue clock. Cannot lock electronics.
Channel 7 on the V124. Rob Mitchnoff and Angelika set up the clocks.
Keep present store until after 10:00 AM so Waldo and Rob can look at it.
Background rates in PHENIX are high. Will need Gap Cleaning, at least in
Gold. Acceleration RF cavities raised from 75 to 100 kV. It improved
luminosity.
Yun will work on decoupling during the evening shift today.
Present store will be kept until 11:00 AM.
STAR Beam-Beam counters like right. ZDCs look got, but the TAC is broken.
Finish rebucketing during the day using the rest of this store. Injection
drift will get worked on for the next fill. And gap cleaning will get done
last.
Angelika will come in to do the Blue collimators later today.
Auto-cogging not working.
Experiments will get the night shift again tonight.
19 Nov 2007: Yesterday, tune feedback and ramp replay go
going in Blue, but not yet in Yellow. 37x37 bunches at 90% ramp
efficiency. Yellow tune path was altered to keep away from the 60 Hz line
by Steve T. This was put into the design edit\or for the machine for the
tune we want, instead of the model tune. There was a major confusion about
what was the tune of the machine and what to use for tune feedback. GregM
will address this problem.
The vacuum at L05 and L07 were fixed, with vacuum better than ever.
Plan for today is to get the experiments magnets on. Angelika needs some
time to set up the collimators. There are some dead loss monitors that
need fixing. If there is nothing simple in the
service building, there will need to be an access to fix the loss monitors.
Not Yet Ready to start giving the owl shift to experiments.
Angelika needs about 4 hours of owl shift tonight to work on collimator set-up.
PHENIX also would like a 15 minute access to move in the
second carriage and check for cleanliness for magnet.
STAR would take advantage of a 1 hour access to change some trigger cables to
correct timing.
Jonathan also needs a 1 hour access to tie in the PHENIX Iris Scanner to make it
functional.
RF needs about 1 hour with beam in each ring to do the rebucketing. Also
needs to do phasing at flattop. Happy to do it at 6 bunches.
There will be an access today from 13:00 - 14:00 to allow for this work.
18 Nov 2007: Increasing bunches through 12x12, 28x28 to 37x37 giving ZDC rates of ~15kHz
at both STAR and PHENIX and beam currents of 7E10 deuterons and 6E8 gold.
Single beam (Blue) running with 60 and 110 bunches. Today go to colliding
56x56 first. Then work on fixing the beam radius. By midnight tonight, we
could have collisions of 56x56. The experiments can start timing in their
detectors.
Continued vacuum work in AGS, this time at L07 where replacing a flange on a
blank-off valve seems to have stopped the leak for the moment. Ion pump at
L07 tripped off. Bad vacuum was responsible for RF tripping off.
Feedback and replay working for Blue. Ramp efficiency at 90% for bleu
with ARTUS off and feedback on, somewhat less for Yellow. No feedback yet
for Yellow. There appears to be an overlap with a 60 Hz resonance and the
Yellow tune. The Yellow tune needs to be changed in the Ramp, although
Larry offered to put it in by hand so that work on Yellow could go forward.
Steve Tepekian, Jim Cameron, Ron Schroeder will work on getting feedback working
for Yellow by getting BBQ working, improving the tunes in Yellow to avoid 60Hz
overlap.
Angelika needs to work on the collimators.
Cavity conditioning, 48 hour estimate, was overly conservative. We are
in a position to start rebucketing tomorrow.
17 Nov 2007: Beams collide with six bunches! Both
STAR and PHENIX see ~800Hz collision rate. Efficiencies for Blue and
Yellow were both at 85%. Will try increasing to 12 bunches today.
But vacuum problems in the AGS at L05
have reoccurred. It appears that both the seal and the weld
between the flange and the vacuum chamber were damaged by high losses in
the gold beam.
Continued problems with tune feedback. Turning on BBQ seems to
make the tune stable. Will continue work on tune feedback tonight
at 8:00 PM when Yun comes back. Larry will be available, either
from home or will come in.
16 Nov 2007: Vacuum problem in L05. Beam is not getting through, and vacuum pumps
tripped off. J10 bump not operating for the deuteron cycle. The puts
the beam loss on the inside of the ring. This damaged the kapton (burned)
and broke the vacuum. It was sprayed with sealant, which allowed the
vacuum pumps to be turned on again. The J10 bump running full current is
not enough to keep from losses.
Carl finished his PS work.
Freddy put in
new RF scaling factor for the RF code. Blue/Yellow synchronization has been
fixed.
Vadim found a vertical corrector with the wrong polarity. It has been
put on STBY and will be fixed today. It may be responsible for the trouble
making orbit corrections. He also found a bad BPM and put it off.
Tune/Coupling feedback loops need to be closed. Closing the loops now appear to
increase the coupling rather than decrease it. So until Yun can work on
the tune/coupling feedback, we will work just on improving the ramp and tune.
BBQ locks fine on Blue. Getting BBQ to work on Yellow would be good.
15 Nov 2007: Problems getting the new slow factor 1 ramp going in Yellow (dAu82).
The new ramp made things worse.
STAR completed their magnet tests for their
white sheeting.
Transmission is at 82%Blue and 55% Yellow.
Quench ground
current monitoring connector fixed.
Carl Schultheiss needs 1 hour to install
new software for the MMPS
There is a bug that allows the correction to
be applied more than once. Until this software bug gets fixed, better
communication will help to avoid this in the future.
14 Nov 2007: Need 1 hour to adjust tune and coupling correction. Currently it
is being tuned by hand. BBQ system is not tracking the tune.
Did not agree with ARTUS. BBQ BTF thinks the tune is one thing,
and BBQ LOCK thinks it is something else. Larry Hoff and Al Dellapena need to get together and agree on what the problem is, and the
approach to a solution.
Need to go to slow factor 1. Don Bruno
needs 3-4 ramps, then Carl Schultheiss needs to replace the connector to
the Yellow Quad quench ground current that is reporting a spurios
current. Then he will need 1-2 ramps to see that it is working
well. This requires 2 hours.
Chipmunk failure has put in all
beamstops all the way back to Tandem.
Both rings are now ramping
together. Blue transmission 74%, Yellow 55%.
Yellow orbit
correction on the ramp not working well. Todd should measure the
lattice. Gamma_T setup needs work. Yellow frequency not
synchronized with Blue. RF group can look at this behind ramps.
13 Nov 2007: Working on blue ramp. Try to do ramp
development with tune feedback. Need to have Larry there to do this work.
Angelika will correct the orbit in Blue, for however long this store lasts.
(Oops, store was accidentally tossed.) At the end of this store, Controls
will go in to replace the DAC board. There is a water leak in the 912
cooling tower for Booster RF. There will be a 2 hour down time during
which the cooling water leak will be repaired, and the DAC board will be
replaced with a spare.
Yellow PS work, waiting for ramp to be fixed.
Steve will make the yellow ramp at slow factor of 1 for 90% of the ramp, and
slow factor 2 for the last 10%. Steve will create a new ramp, dAu82, that
extends the beta squeeze on Yellow. Blue ramp will be unchanged
Steve estimates that this work will take ~1 day. But we still have large
radial excursions, as we go up the ramp. It is better to fix the ramp than
to put in trim magnet settings which would be very big. So Steve will
modify the ramp instead. The Yellow ramp that is currently installed has
an 8 mm orbit excursion at transition.
10 Nov 2007: Gold injected into Yellow
Single bunch of deuterons circulated in Blue from 2:00 AM til end of shift.
Power supply work on Yellow Gamma_Ts to start at 10:00 AM.
There will be an access of ~2 hours to allow for Yellow PS work. STAR and
PHENIX have been notified.
09 Nov 2007: RHIC is off this morning to allow for ACS to work on PHENIX flammable gas alarms in the experimental hall. Work should take from 7AM to 11AM. It should be possible to complete STAR magnet testing at the same time. Yellow PS should be tested to Injection energy today. Shutoff tests from injection will be run today, but shutoff from high current will fire the Blue abort kicker due to cross-talk. Bringing Yellow to a full hysteresis ramp should happen sometime on Sunday.
08 Nov 2007: Work on A5 to narrow the pulse. Work completed, but the A5 kicker pulse has a bad undershoot. Now injection shows all 8 deuteron pulses being transferred. However, the F3 kicker showed problems, intermittent connections. Work on A5 took less time than anticipated, which was good because the F3 kicker in the Booster started misbehaving and the Booster had to be LOTOed to allow the F3 to be repaired. That work went well eventually. Cooling tower 6 and the fan causing trips of the AtR due to overheating. Todd got the WCM working, and the deuteron beam current calibrated.
07 November 2007: Work continues on the Blue MMPS and correctors. The plan is to bring the Blue MMPS to injection energy by 7PM tonight, and then work on the full hysteresis by midnight. Assuming this goes well, there will be an attempt at that point to capture beam in RHIC and circulate it to check for obstructions.
01 November 2007: There were problems yesterday setting up for the power outage.
so attempts to recover deuterons were abandoned until after the LIPA shutdown. Many of the difficulties experienced today could have been reduced with the help of better checklists (for the sweeping of the TtB) and improved procedures (such as keeping the TtB running when there is a UPS failure in the EBIS/LINAC complex.)
22 October 2997: At the Scheduling Meeting today it was decided that the Q7 polarity would be reversed to allow the Intra Beam Scattering (IBS) Suppression lattice to be used during the d-Au running. This means there will be a 8-12 hour job (LOTO, cable splicing, documentation) to accomplish the switch. There will be an additional vacuum break for the 90MHz RF cavity. A similar interruption will be required at the switch to polarized proton running to undo these changes. It was viewed as a net positive to make the changes. ... We discussed the move for the ZDCs to account for the different crossing angle used in d-Au running, and the need to undo this change as well before starting pp. Ed O'Brien said the Shower Max detectors should move with the ZDCs. Angelika recorded the move as 1 cm towards the yellow ring. ... There was a discussion of when the experiment magnets should be turned on. It was agreed that it should be as late as possible to save costs. (Angelika remembered that the orbit tune required to compensate for the magnets took only a short time.) So experiments should plan of bringing their magnets on ~1 week after RHIC is on, or ~ Nov 12 .... Both experiments requested time to flip their magnet polarity one time during the d-Au running and again during the pp running. ... Tonight deuteron development starts. Tandem has had problems with the deuteron source. ... The LIPA work on the Booster feed is scheduled for 1 November 2007. ... There will be a power outrage on 24 October startingat 7:00 at 12 OClock, until 14:00, and also beginning at 10:30 and ending at 16:30 at the 10 OClock Hall. ... The Second RHIC Dry Run will be Oct 29 - 31.
11 October 2007: Tom Nehring sent out a notification of power outages during the next week. Joe Tuozzolo sent out a schedule of Liquid Helium delivery. Yousef Makdisi reported on the installation of new replacement silicon detectors for the JET Polarimeter.
10 October 2007: RHIC 08 Dry Run continues, on to Day 2. A decision has been made to schedule the LIPA Transformer work for Thursday 1 Nov 2007, with Tuesday 30 Oct as a secondary choice. The LIPA work requires the Booster Main Magnet power to be off.
08 October 2007: Fast Extracted Beam to NSRL.
27 September 2007: Kevin Brown called to ask if Betsy Sutherland would be using Fast Extracted Beam in her next run. The fast kicker is being repaired in the Booster. Ray reported that kicker repairs were complete, and he was in the middle of testing. He will report the results for those tests. Kevin was also concerned about the status of the BPMs at E3, E5 and E8.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email:sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |